OpenStudy (anonymous):

six friends are going to a play. The theater offers a $4 discount toward each $15 ticket. a. Find the total amount of money the friends will spend on the play

OpenStudy (anonymous):

There are a couple of ways you could solve this... let's try this way first: If each ticket costs $15, but the theater is giving a $4 discount for each ticket, then how much does each person have to pay? They don't have to pay the full $15, right? So how much do they pay after they get the discount?

OpenStudy (anonymous):

15 x 6 = 90 6 x 4 = 24 90 - 24 = 66$ in all. They each pay 11$

OpenStudy (anonymous):

The question asks for the total amount of money, so I wouldn't answer $11... that's the per-person price after discount.

OpenStudy (anonymous):

I meant 66$ in all.

OpenStudy (anonymous):

But you do have the total there... I just didn't want @foreveryoung615 to miss it.

OpenStudy (anonymous):

Yes, @foreveryoung615, @ValentinaT showed the other way to solve. One way is find the price per person after discount... 15 - 4 = 11 and then multiply by the number of people to get the total: 11 x 6 = 66 @ValentinaT showed that you get the same answer if you find the total full price first, by multiplying the full ticket price by # of people: 15 x 6 = 90 and then multiplying the discount by # people: 6 x 4 = 24 and then subtracting the total discount from the total full price: 90 - 24 = 66